# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

### v1

@@ 3.2.3. Self-Refinement
Based on these aspects, we prompt the LLM to give a numerical score out of 5 and select the questions that have a full score of 25. This filter-feedback-refine process is repeated once more to ensure better quality of questions. Refer to Appendix C.2 for the detailed prompts.
Table 1. Number of questions and proportional decrease in each benchmarking task before and after Self-Refinement.
Task Num. (Before) Num. (After) % Decrease
AttackSeq-Tactic 2,158 1,697 -21.36
AttackSeq-Technique 2,937 1,917 -34.73
AttackSeq-Procedure 4,642 2,635 -43.24
3.2.4. Dataset Statistics
The number of generated questions for each benchmark task, along with the corresponding percentage decrease, is shown in Table 1. More details of the dataset, such as the distribution of ATT&CK tactics and techniques associated with the questions, is in Appendix A.1.
3.3. Dataset Evaluation
3.3.1. Human Evaluation
We ask two postgraduate students majoring in computer science, who are knowledgeable about cybersecurity and the MITRE ATT&CK framework, to assess the quality of questions in AttackSeqBench. We randomly sample 35 questions from each of the three benchmarking tasks, including the AttackSeq-Procedure-No subtask, resulting in a total of 140 questions for manual evaluation. Each annotator evaluates the six aspects described in Section 3.2.3 using a 5-point Likert scale, where higher scores indicate better alignment (detailed definitions of each aspect are in Appendix A.2). Based on Table 8, the average accuracy score of annotators over the three tasks is 0.625, suggesting that questions in AttackSeqBench are notably challenging for domain experts. Furthermore, the human evaluation results indicate that the generated questions closely satisfy with each aspect of the evaluation criteria, with an average score of 4.620/5 across the six aspects. In particular, the Logical and Relevance aspects suggests that LLMs are able to construct questions that closely align with attack sequences described in CTI reports.
3.3.2. Automatic Evaluation
To alleviate the laborious task of manual human annotation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge frameworks in aligning with human preferences within domain-specific fields, including cybersecurity (Xu et al., 2024b). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated questions. Similar to Section 3.3.1, we carefully design prompts for each of the six aspects using a 5-point Likert scale (refer to Appendix C.3 for the detailed prompt). More specifically, we instruct GPT-4o to refer to the evaluation criterion’s definition, along with description of the correct answer from the ATT&CK KB, and rate the question out of 5 for each aspect. Based on Table 9, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences. While the automatic evaluation scores are lower than the human evaluation scores, this implies that it is more challenging for LLMs to answer the questions correctly as compared to domain experts.
4. Experiment
Figure 3. The illustration of three benchmark settings. (a): Regular setting, (b): Zero-shot setting, and (c): RAG-empowered setting.
4.1. Benchmark Settings
As illustrated in Figure 3, we elaborate on the three proposed benchmark settings, Regular, Zero-Shot and RAG-Empowered. The detailed prompts for each setting is in Appendix C.4.
Regular Setting. Inspired by existing works (Ma et al., 2023; Jin et al., 2024), we propose this benchmark setting to assess the impact of contextual information in analyzing attack sequences. We adopt an ICL-based approach by carefully designing a well-constructed prompt using CoT. Within the prompt, we first instruct the LLM to analyze the attack sequence described in the masked CTI outline associated with the question, and subsequently deduce the correct answer based on the given attack sequence.
Zero-Shot Setting. Aside from excelling in semantic understanding, previous works have also demonstrated LLMs’ impressive zero-shot reasoning ability across various downstream tasks (Hou et al., 2024; Kojima et al., 2022). To evaluate the LLMs’ capabilities in zero-shot attack sequence analysis, we omit the masked CTI outline from the prompt in the Regular setting. Instead, we instruct the LLM to select the most plausible answer choice based on the attack sequence described in the question’s body.
RAG-Empowered Setting. Retrieval Augmented Generation (RAG) has gained traction due to its ability to enhance LLMs’ reliability and mitigate hallucinations (Zhang et al., 2025a). We introduce this benchmark setting to assess the impact of providing the relevant ATT&CK TTPs in the question’s body as contextual information, by leveraging a retrieval-based approach and the ATT&CK KB. We first crawl the description and the example procedures of each TTP from the Enterprise ATT&CK Matrix  [^1] https://attack.mitre.org/matrices/enterprise/. Next, we split and embed the text chunks alongside their associated ATT&CK tactic and technique as metadata. Afterwards, we retrieve relevant chunks based on the question’s body and extract the associated ATT&CK TTPs from the metadata. Lastly, we instruct the LLM to utilize the retrieved TTPs and deduce the correct answer. By doing so, we enable the LLM to focus on determining the correct attack sequence rather than interpreting the TTPs in the question.
Table 2. Performance (Accuracy) comparison of various open-source and proprietary LLMs (both fast-thinking and slow-thinking reasoning LLMs) under three settings (i.e., Regular, Zero-Shot, RAG) on three datasets (i.e., Tactic, Technique, Procedure). For each column, bold number indicates the best performance, and underlined number corresponds to the second best performance.
LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
Regular Zero-Shot RAG Regular Zero-Shot RAG Regular Zero-Shot RAG
Fast-thinking LLMs
[…]
Qwen-2.5-7B-Instruct 0.5121 0.4903 0.4761 0.6693 0.6568 0.6067 0.6584 0.5184 0.4941
Llama-3.1-8B-Instruct 0.4744 0.5085 0.4926 0.6260 0.6333 0.5827 0.6577 0.5328 0.5230
ChatGLM-4-9B-Chat 0.4885 0.4979 0.5009 0.6275 0.6109 0.6030 0.641 0.5408 0.5131
Llama-3.3-70B-Instruct 0.6588 0.5551 0.5681 0.7058 0.6797 0.7037 0.6903 0.5469 0.5279
Qwen-2.5-72B-Instruct 0.5793 0.5863 0.5657 0.5430 0.7162 0.6959 0.7188 0.6285 0.6030
GPT-4o-mini 0.6517 0.6005 0.5692 0.7387 0.7058 0.7021 0.6968 0.5491 0.5340
GPT-4o 0.6093 0.5740 0.5787 0.6755 0.6995 0.7188 0.7359 0.6755 0.6353
Slow-thinking Reasoning LLMs
[…]
DeepSeek-R1-Distill-Qwen-32B 0.5421 0.5698 0.5504 0.5879 0.5816 0.5816 0.6945 0.6258 0.5852
QWQ-32B-Preview 0.5345 0.3377 0.4638 0.5112 0.3918 0.5342 0.7036 0.5696 0.5457
GPT-o3-mini 0.4643 0.5445 0.5215 0.5373 0.5915 0.5822 0.6854 0.6877 0.6459
4.2. Experimental Setting
4.2.1. Implementation Details
To demonstrate the robustness of AttackSeqBench, we evaluate eight fast-thinking LLMs: Mistral-7B-Instruct-v0.3, Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct, ChatGLM-4-9B-Chat, Llama-3.3-70B-Instruct, Qwen-2.5-72B-Instruct, GPT-4o-mini and GPT-4o. Additionally, we assess four slow-thinking reasoning LLMs: DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-32B, QWQ-32B-Preview and GPT-o3-mini. For the RAG-empowered benchmark setting, we utilize a hybrid retriever due to its considerable performance in existing work (Zhang et al., 2024b). We combine Okapi BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on BGE-EN-ICL (Li et al., 2024) (shortened as BGE) and Chroma DB  [^2] https://www.trychroma.com/, along with a re-ranker. The detailed implementation is in Appendix B.1.
4.2.2. Evaluation Metrics
We evaluate the Accuracy (abbreviated as Acc in the following sections) for all three tasks in AttackSeqBench (Liu et al., 2024; Zhang et al., 2024b; Alam et al., 2024). Specifically, given the total number of questions $M$ and the number of questions $n$ with the correct answer choice chosen by the LLM, the accuracy can be calculated as $Acc=n/M$.
4.3. Performance Comparison
We conduct comprehensive analysis using eight fast-thinking LLMs and four slow-thinking reasoning LLMs in Table 2 and elaborate on three key findings: (1) None of the LLMs clearly outperforms other LLMs across all the tasks. While GPT-4o demonstrate strong performance in AttackSeq-Technique and AttackSeq-Procedure, its performance falls behind Llama-3.3-70B-Instruct and GPT-4o-mini in AttackSeq-Tactic. This implies the complexity and the diversity of our three sub-tasks to some extent. (2) Among these three tasks, we observe that LLMs collectively demonstrate the lowest overall performance for the AttackSeq-Tactic task. This implies LLMs struggle to determine the high-level objectives in an attack sequence, even with the CTI outline given in the context. This may arise due to the abstract nature of the descriptions of tactics which causes the LLM to misinterpret the correct tactic. (3) The AttackSeq-Procedure task yields the highest performance results across all the LLMs. This suggests that LLMs are able to leverage its procedural knowledge to deduce the plausibility of a group of procedures in an attack sequence. However, this may also be attributed by the LLMs’ tendency to favor ”yes” in boolean questions when given insufficient context (Wen et al., 2024), which causes high accuracy for AttackSeq-Procedure-Yes. We explore this phenomenon further in Section 4.4.1.
4.4. Model Analysis
Table 3. Performance comparison of four representative LLMs between AttackSeq-Procedure-Yes and AttackSeq-Procedure-No sub-tasks in each benchmark setting. For each column, the bold value indicates the best performance.
LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
Regular Zero-Shot RAG Regular Zero-Shot RAG
Llama-3.1-8B 0.7899 0.9518 0.9460 0.5432 0.1700 0.1565
GPT-4o 0.7146 0.9403 0.9542 0.7542 0.4462 0.3591
R1 (Llama-8B) 0.8332 0.9526 0.9477 0.4341 0.1161 0.1062
GPT-o3-mini 0.5789 0.7760 0.8242 0.7776 0.6112 0.4915
4.4.1. Impact of Contextual Information
To comprehensively investigate the impact of the contextual knowledge in the given CTI outlines for analyzing attack sequences in LLMs, we compare the performance of LLMs between the Regular and Zero-Shot benchmark settings. Based on in Table 2, we observe that both fast-thinking and slow-thinking reasoning LLM’s accuracy generally increases across each task in the  Regular setting. Among the fast-thinking LLMs, the performance of models with smaller number of parameters, e.g., Llama-3.1-8B-Instruct (shortened as Llama-3.1-8B), have minimal or no improvement for AttackSeq-Tactic and Attack-Technique tasks. This indicates that the baseline LLMs with smaller parameters may not fully leverage the capabilities of ICL within these benchmarking tasks. Similarly for slow-thinking reasoning models, DeepSeek-R1-Distill-Llama-8B (shortened as R1 (Llama-8B)) has no improvement in the Regular setting as compared to the Zero-Shot setting. On the other hand, Llama-3.3-70B-Instruct achieves comparable performance against the closed source models, i.e., GPT-4o and GPT-4o-mini. Among the three tasks, AttackSeq-Procedure has the highest improvement in LLMs from the Zero-Shot setting to the Regular setting. Based on Table 3, we observe a significant increase in the AttackSeq-Procedure-No sub-task. This suggests that the contextual information found in the masked CTI outline is essential in justifying procedures that are unlikely to happen.
Table 4. Performance comparison between three embedding models. Eeb. M. refers to the embedding model used, while Tac, Tec and Pro refers to the three tasks in AttackSeqBench (i.e., Tactic, Technique, and Procedures) respectively. For each row, the bold value indicates the best performance.
#Eeb. M. Task Llama-3.1-8B GPT-4o R1 (Llama-8B) GPT o3-mini
BGE Tac 0.4926 0.5787 0.4532 0.5215
Tec 0.5827 0.7188 0.5138 0.5822
Pro 0.5230 0.6353 0.4968 0.6459
OpenAI Tac 0.4915 0.4814 0.4632 0.5333
Tec 0.5790 0.7178 0.5039 0.5962
Pro 0.5195 0.6330 0.4979 0.6353
ATT&CK- BERT Tac 0.4985 0.5622 0.4590 0.5292
Tec 0.5790 0.7100 0.5018 0.5874
Pro 0.5245 0.6209 0.4953 0.6398
4.4.2. Effectiveness of RAG Strategies
We analyze the capabilities of the RAG-empowered setting based on the results of AttackSeqBench in Table 2. Within the set of fast-thinking LLMs, GPT-4o achieves exceptional performance when leveraging RAG in AttackSeq-Tactic and AttackSeq-Technique tasks. Moreover, the LLMs exhibit the best performance in the AttackSeq-Technique among the three tasks in Table 2. This indicates that LLMs are able to leverage the retrieved TTPs to determine the correct technique. However, the RAG-empowered setting does not consistently yield effective results, as models that have less than or equal to 9B parameters have no improvement as compared to the Zero-Shot setting. This suggests that the information retrieved, i.e., the semantically similar ATT&CK TTPs based on the question’s body, is insufficient for smaller models. Furthermore, we evaluate the performance of four representative LLMs using two additional embedding models in Table 4. The two models are: OpenAI’s text-embedding-3-large (OpenAI, 2024c); ATT&CK BERT (Abdeen et al., 2023) which is an open-source embedding model fine-tuned for cybersecurity. Based on Table 4, BGE acheives the best performance across the three benchmarking tasks.
Figure 4. Performance and model size comparison of nine representative LLMs in the AttackSeq-Tactic task. The size of bubble corresponds to inference time, where horizontal lines denote fast-thinking LLMs and vertical lines indicate slow-thinking reasoning LLMs.
4.4.3. Computational Complexity Analysis
As illustrated in Figure 4, we extensively compare the performance, model size and inference cost of nine open-source LLMs in the AttackSeq-Tactic task. We have the following observations: (1) In general, these LLMs exhibit the performance trend in sequential attack pattern understanding which aligns with the scaling law (Kaplan et al., 2020). (2) Comparing the performance of LLMs with different thinking types, we observe that slow-thinking LLMs cannot significantly and consistently outperform the other. One possible reason is that given the question and masked CTI outlines, fast-thinking LLMs can already deduce the correct answer without any elaborate reasoning. (3) Regarding the inference cost, we notice that slow-thinking LLMs do not incur more reasoning costs compared to fast-thinking LLMs. This prompts an intuitive insight: the small-scale slow-thinking reasoning LLM fine-tuned using AttackSeqBench may enhance its capability in attack sequence analysis significantly while maintaining the lower reasoning cost.
4.4.4. Sampling Parameters Analysis
We assess the impact of two sampling parameters on the performance of LLMs in AttackSeq-Tactic under the Regular setting, namely temperature and maximum output tokens as illustrated in Figure 5. We observe that increasing the temperature at different three intervals has minimal improvements for Llama-3.3-70B-Instruct, while resulting in negative performance for the remaining three LLMs. This indicates that the most optimal performance occurs when the temperature is kept at 0.0. On the other hand, we observe a considerable improvement when increasing the number of tokens for slow-thinking reasoning models. The accuracy of Deepseek-R1-Distill-Llama-8B improved by 8.46% as the maximum output tokens increased from 1024 to 4096, while Deepseek-R1-Distill-Qwen-32B achieve a 24.9% improvement for the same increment. However, further increasing the token limit to 8192 yields diminishing returns for both models. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in slow-thinking models while considering the associated costs incurred (Wang et al., 2024).
Figure 5. Parameter analysis on temperature (a) and maximum output tokens (b) in AttackSeq-Tactic task using the Regular setting, we perform analysis on four representative LLMs using different colors to represent them.
Figure 6. The performance comparison of four representative LLMs on each tactic in each benchmarking task.
Table 5. Reasoning path comparison between slow-thinking reasoning and fast-thinking LLMs in Zero-Shot setting based on the AttackSeq-Procedure task in Zero-Shot setting.
Question: After delivering spearphishing emails with spyware disguised as a PDF file to gain initial access, is it likely the authors of Hacking Team’s spyware established communication with the identified C&C servers before using VMProtect for obfuscation to evade detection? Answer Choices: A) Yes  B) No (Correct Answer)
DeepSeek-R1-Distill-Qwen-32B Qwen-2.5-72B-Instruct
First, spearphishing emails are emails that look legitimate, which are actually malicious. In this case, the PDF is a disguise for spyware. So the first step is the delivery of the malicious payload via email. (Knowledge Retrieval) Once the victim opens the PDF, the spyware is executed. The next step in an attack chain is usually establishing communication with a command and control (C&C) server. Now, the question mentions VMProtect being used for obfuscation. However, sometimes obfuscation can be applied after deployment to maintain persistence or evade detection during operation. (Breakdown of Attack Sequence) But in the context of the attack sequence described, the spyware has already been delivered and executed. The next logical step would be for the spyware to reach out to the C&C server to report back and receive instructions. (Self-Correction) If the attackers were to use VMProtect at this point, it might be to re-obfuscate the spyware or to protect it further, but that seems a bit out of order. (Reasoning using Attack Sequence) So the answer would be No. (Correct Conclusion) The sequence includes: Delivering spearphishing emails with spyware disguised as a PDF file to gain initial access. Establishing communication with C&C (Command and Control) servers. Using VMProtect for obfuscation to evade detection. (Breakdown of Attack Sequence) The spearphishing emails with a PDF file are used to gain initial access to the target system. (T1566.002: Spearphishing Attachment). Once the initial access is gained, the malware typically establishes communication with the C&C servers to receive further instructions or exfiltrate data. (T1071: Application Layer Protocol). VMProtect is a tool used to obfuscate the malware to evade detection by security software. (T1027: Obfuscated Files or Information). (Knowledge Extraction) The logical sequence of events would be: 1. **Initial Access:**, 2. **C&C Communication:**, 3. **Obfuscation:**. (Incorrect Attack Sequence Order) Based on the logical sequence of TTPs, it is likely that the authors of Hacking Team’s spyware established communication with the C&C servers before using VMProtect for obfuscation to evade detection. Final Answer: Yes (Incorrect Conclusion)
4.5. In-depth Analysis
4.5.1. Impact of TTP Temporal Position
To better understand the relationship between LLMs and attack sequences, we conduct fine-grained analysis on the performance of four representative LLMs under the Regular benchmark setting and across the ATT&CK tactics (refer to Figure 6). Additionally, we identify four categories to group the tactics into overarching attack phases as shown in Figure 6: Initial Intrusion, Exploitation, Stealth Expansion, Objective Orchestration. Our findings are: (1) Based on the standard deviation range across various tactics in each LLM, we observe that the range is the largest for AttackSeq-Tactic, indicating that the LLM’s performance largely varies for this task across different tactics. In contrast, the SD range is the smallest for AttackSeq-Procedure which suggests more stability in the LLMs’ performance. This highlights the challenge for LLMs to determine the correct tactics in attack sequences, as compared to procedural knowledge needed to correctly determine techniques and procedures. (2) GPT-4o consistently outperforms other models across most tasks, particularly in the in the Exploitation and Stealth Expansion phases. On the other hand, Llama-3.1-8B-Instruct and GPT o3-mini achieve comparable performance against GPT-4o in Initial Intrusion and Objective Orchestration phases, especially at the start and end of the sequence (i.e., Reconnaissance and Impact respectively). This inconsistency suggests that TTPs within certain ATT&CK tactics are more challenging for the LLM to interpret, likely due to the rarely observed tactics in real-world CTI reports. More detailed distribution of the performance can be found in Section B.2 of the Appendix.
4.5.2. Inference Path Analysis on Slow-thinking Reasoning LLMs
[…]

### v2

@@ A.2 Dataset Evaluation Criteria
To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table LABEL:tab.annotation_instructions), where each score corresponds to a different level of the given aspect. Then we instruct three cybersecurity experts and LLM to provide the score of each aspect to achieve the human evaluation and the automatic evaluation, respectively. The detailed results are shown in Table 1. While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most Q&A pairs in the dataset satisfy the requirements of all aspects. This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity. Note that for the AttackSeq-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeq-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
A.3 Benchmark Tasks
Inspired by existing LLM benchmarks in the general domain (Hendrycks et al., 2021a; Rein et al., 2023; Zhang et al., 2024b), we propose three tasks in the form of Multiple-Choice Questions and Yes-No Questions to evaluate the reasoning capabilities of LLMs in inferring TTPs present in attack sequences, where each task reflects a distinct aspect of adversarial behaviors.
Table 5: Annotation instructions for each evaluation aspect.
Aspects Instructions
Answerability Score 1: The correct answer is not supported by the CTI outline. The information is either missing or contradicts the correct answer. Without the masked tactic paragraph, it is impossible to deduce the correct answer.
Score 2: Some evidence in the CTI outline loosely supports the correct answer, but it does not clearly stand out as the best choice. Removing the masked tactic paragraph makes it highly difficult to deduce the answer, even when referring to the MITRE ATT&CK KB.
Score 3: The correct answer has partial support in the CTI outline but is not explicitly stated. After removing the masked tactic paragraph, it is possible but challenging to infer the correct answer using the remaining information and MITRE ATT&CK KB.
Score 4: The correct answer is well-supported by the CTI outline and is the most reasonable choice based on the provided information. If the masked tactic paragraph is removed, the answer remains largely deducible using remaining information, and MITRE ATT&CK KB.
Score 5: The correct answer is directly supported by the CTI outline and is unambiguously the best choice. Even if the masked tactic paragraph is removed, the answer remains easily deducible based on the remaining CTI outline and MITRE ATT&CK KB.
Clarity Score 1: The question is highly ambiguous, imprecise, or contains vague phrasing. It may directly state the correct answer, making inference unnecessary.
Score 2: The question is somewhat unclear or contains minor ambiguities. It may hint too strongly at the correct answer, reducing the need for inference.
Score 3: The question is fairly clear, but minor ambiguities exist. It does not directly state the correct answer, but slight rewording could improve precision.
Score 4: The question is mostly clear and unambiguous. It requires inference and does not directly reveal the correct answer.
Score 5: The question is precise, completely unambiguous, and free of vague phrasing. The correct answer is never directly mentioned, ensuring inference is required.
Logical Score 1: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline.
Score 2: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics.
Score 3: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly.
Score 4: The question follows the sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner.
Score 5: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer.
Relevance Score 1: The question is completely unrelated to the CTI outline.
Score 2: The question has only slight relevance to the CTI outline but is mostly off-topic.
Score 3: The question is somewhat related to the CTI outline but could be refined to better fit the content.
Score 4: The question is directly related to the CTI outline, with minor room for improvement.
Score 5: The question fully aligns with the CTI outline and is highly relevant to the content.
Consistency Score 1: The question contradicts the TTP description or is entirely misaligned with the provided details.
Consistency Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
Score 3: The question mostly aligns with the TTP description but contains minor inconsistencies.
Score 4: The question is highly consistent with the TTP description, with only minor areas for improvement.
Score 5: The question fully aligns with the TTP description, with no inconsistencies or contradictions.
Answer Consistency Score 1: The correct answer does not fully resolve the question, leaving contradictions or gaps.
Score 2: The correct answer provides some resolution, but contradictions or inconsistencies remain.
Score 3: The correct answer is mostly consistent, but minor contradictions exist.
Score 4: The correct answer fully resolves the question with minimal inconsistencies.
Score 5: The correct answer completely and unambiguously answers the question, with no contradictions or inconsistencies.
AttackSeq-Tactic. This task evaluates the LLMs’ capability to infer a tactic $t_{k}\in T$. Given a question $Q$ that corresponds to tactic $t_{k}$ and four shuffled candidate tactics $C_{Tac}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct tactic $c_{l}\in C_{Tac}$.
AttackSeq-Technique. This task assesses the LLMs’ capability to infer a technique $e_{j,k}\in E(t_{k})$. Given a question $Q$ that corresponds to $e_{j,k}$ and four shuffled candidate techniques $C_{Tec}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct technique $c_{l}\in C_{Tec}$.
AttackSeq-Procedure. This task challenges the LLMs’ capability to determine the likelihood of procedures $p_{m,j,k}\in P(e_{j,k})$ in an attack sequence. Given a question $Q$ and two candidate choices $C_{Pro}=\{yes,no\}$, the LLM will be instructed to determine if the procedure $p_{m,j,k}$ is likely to occur in the given attack sequence $S$.
We further divide AttackSeq-Procedure into two sub-tasks, namely AttackSeq-Procedure-Yes and AttackSeq-Procedure-No, based on the ground truth of the boolean question. This explores the LLMs’ ability in determining misleading procedures that are unlikely to occur in an attack sequence.
A.4 Baselines
To demonstrate the effectiveness and robustness of our proposed AttackSeqBench, we evaluate seven large language models, five large reasoning models and four post-training strategies across three tasks involving different levels of data and three benchmark settings with varying context completeness. We leverage vLLM (Kwon et al., 2023) to run all the open-source LLMs locally with two Nvidia H100 GPUs. For the colsed-source LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to conduct inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below is the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
Large Language Models:
  • LLaMa-3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • ChatGLM-4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen-2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
  • LLaMa-3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
  • GPT-4o (OpenAI, 2024a) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing.
Large Reasoning Models:
  • DeepSeek-R1-Distill-Llama-8B (R1 (Llama-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the Llama-3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
  • QWQ-32B-Preview (QWQ-32B) (Team, 2024) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
  • GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities. It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
Post-Training Strategies:
  • Supervised Fine-tuning (SFT) (Zhang et al., 2023) is a critical process for adapting pre-trained LLMs to specific tasks by training them on a task-specific dataset with labeled examples.
  • Reasoning Distillation (RD) (Huang et al., 2024) RD is a widely adopted approach for enhancing LLM reasoning, which collects reasoning samples with self-reflection from existing LRMs and distills them to guide LLMs in acquiring long-thought capabilities.
  • Reinforcement Learning from Internal Feedback (RLIF) (Zhao et al., 2025) replaces the external rewards in Group Relative Policy Optimization (GRPO) with LLMs’ self-certainty, enabling unsupervised learning from intrinsic signals without relying on external rewards.
  • Reinforcement Learning with Verifiable Rewards (RLVR)  (DeepSeek-AI, 2025) leverages rule-based verification functions to provide reward signals for tasks with clear correctness criteria, enabling the optimization of LLMs while avoiding the complexities and potential pitfalls of reward models within RLHF.
A.5 Post-training Corpus Construction
Considering the post-training strategies, we construct two diverse datasets for SFT and RD, RLIF and RLVR respectively. For the former, we utilize a subset of the Primus-Instruct dataset (Yu et al., 2025). Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document QA, Terraform security mis-configuration repair, and general multi-turn instruction following. To mitigate the inherent bias from linguistic inconsistencies, we filter out non-English samples via the FastText language identification library (Joulin et al., 2016) and manually verify the results, yielding a subset of 710 samples for SFT.
Regarding the latter, we use Primus-Reasoning, a cybersecurity reasoning distillation corpus constructed with DeepSeek-R1 (DeepSeek-AI, 2025) and GPT-o1-preview (OpenAI, 2024a). This dataset includes, but is not limited to, tasks such as Common Weakness Enumeration (CWE) mapping, Common Vulnerabilities and Exposures (CVE) analysis, and multiple-choice questions on general cybersecurity knowledge. Following (Zhang et al., 2025b), we leverage transitional words (i.e., “but”, “however”, “wait”, etc.) as the proxy for inferability, and retain only the 3,890 samples containing at least ten such words when constructing the corpus for RD, RLIF and RLVR.
A.6 Implementation Details
To examine the performance of LLMs on our AttackSeqBench after embedding cybersecurity knowledge, and considering GPU constraints, we evaluate existing post-training strategies on Qwen-2.5-3B (QwenTeam, 2024) and LLaMA-3.1-8B (Grattafiori et al., 2024) under both full-parameter fine-tuning and parameter-efficient fine-tuning paradigms across all benchmark tasks and settings.
Retrieval Augmented Generation (RAG): We first crawl the description and the example procedures of each technique in the Enterprise ATT&CK Matrix v17  [^5] https://attack.mitre.org/versions/v17/matrices/enterprise/. Then, we split the textual data into text chunks and embed the chunks in a vector store (i.e., Chroma DB  [^6] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique. We utilize a hybrid retriever with a re-ranker, by combining combine Okapi BM25 (Robertson & Zaragoza, 2009) and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b). We set the chunk size to 512 and retrieved chunks to 3, and utilize LlamaIndex (Liu, 2022) to implement the retriever. Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023a) within RAG to evaluate their effectiveness in attack sequence analysis.
Supervised Fine-tuning (SFT): We fine-tune the backbone LLM on the first dataset within Appendix A.5 using the LLaMA-Factory (Zheng et al., 2024) framework. Specifically, we deliberately restricted SFT to one epoch with the learning rate of $3\times 10^{-6}$, leveraging DeepSpeed ZeRO Stage-3 with CPU offload for memory efficiency. In our preliminary experiments, extending training process to multiple epochs led to noticeable degradation in the LLMs’ general capabilities outside the cybersecurity domain. This effect can be attributed to “catastrophic forgetting”, where continued exposure to a narrow corpus may overwrite its previously acquired broad knowledge. Thus, a single epoch struck a balance between adapting the LLM to the cybersecurity tasks while preserving its pre-trained general-purpose performance.
Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1  and GPT-o1-preview ), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs. For RD, we fine-tune our backbone LLM on the latter dataset within Appendix A.5 with the same parameter settings in SFT.
Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated. Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where group-normalized rewards reduce variance and stabilize training. We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
Reinforcement Learning with Internal Feedback (RLIF) (Zhao et al., 2025) enables LLMs to optimize policies using intrinsic signals without relying external supervision. In particular, it replaces the external rewards in GRPO with the self-certainty scores, which estimate the LLMs’ confidence from its outputs, to enable fully unsupervised learning while maintaining the stability benefits of GRPO. Here, we conduct RLIF on the same latter dataset within Appendix A.5 using the verl framework, adopting the same training hyperparameters and FSDP set-up as in RLVR.
Figure 8: Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl & Hannon, 2020). Our benchmark emphasizes on the sequences of adversary behaviors described in CTI reports.
A.7 Related Benchmarks Comparison
[…]